OpenStudy (usukidoll):

quick question ... if I have the quadratic u^2-au+b^2 and the quadratic formula is ax^2+bx+c would a = u^2 b=-a and c = b^2?

OpenStudy (anonymous):

Well, it depends, are we saying \(x=u\) or \(x=b\)?

OpenStudy (anonymous):

If \(x=u\), then: \[ u^2-au+b^2 = (1)x^2+(-a)x+(b^2) \]

OpenStudy (usukidoll):

well the standard quadratic formula is \[ax^2+bx+c\]

OpenStudy (anonymous):

\[ a'=1\\b'=-a\\c'=b^2 \]

OpenStudy (usukidoll):

so yes x = u (sorry I know a = 1 I typo) \[u^2-au+b^2\] is what I have , so obviously by the definition \[a = 1, b = -1 ,c =b^2\]

OpenStudy (usukidoll):

ha ha ha my professor made a mistake XD

OpenStudy (usukidoll):

oh sorry b = -a

OpenStudy (usukidoll):

\[a = 1, b = -a ,c =b^2\]

OpenStudy (usukidoll):

so \[b^2-4ac\] \[(-a)^2 - 4(1)(b^2) \rightarrow a^2-4b^2\]

OpenStudy (usukidoll):

\[\frac{a + \sqrt{ a^2-4b^2}}{2}, \frac{a - \sqrt{ a^2-4b^2}}{2}\]

OpenStudy (usukidoll):

so does anyone know an online calculator that can calculate these bad boys to the 4th power?

OpenStudy (anonymous):

Calculate what?

OpenStudy (usukidoll):

I have already calculated them to the 2nd power and already they are big... calculate the last thing I wrote in Latex

OpenStudy (usukidoll):

\[\frac{a + \sqrt{ a^2-4b^2}}{2}, \frac{a - \sqrt{ a^2-4b^2}}{2} \] I need a calculator to calulate these two to the 4th power.

OpenStudy (usukidoll):

I already did\[(\frac{a + \sqrt{ a^2-4b^2}}{2})^2, (\frac{a - \sqrt{ a^2-4b^2}}{2})^2\] byhand and it was crazy

OpenStudy (anonymous):

Oh

OpenStudy (usukidoll):

but I need \[(\frac{a + \sqrt{ a^2-4b^2}}{2})^4, (\frac{a - \sqrt{ a^2-4b^2}}{2})^2\] so is there a code in mathematica or an online calculator that can calculate this monster?

OpenStudy (anonymous):

Just square it a second time

OpenStudy (usukidoll):

oh yeah and distribute like hell... that takes time :(

OpenStudy (usukidoll):

\[\frac{a^2+ 2a\sqrt{a^2-4b^2}+a^2-4b^2}{4}\]

OpenStudy (usukidoll):

\[(\frac{a^2+ 2a\sqrt{a^2-4b^2}+a^2-4b^2}{4})(\frac{a^2+ 2a\sqrt{a^2-4b^2}+a^2-4b^2}{4})\]

OpenStudy (usukidoll):

that's a beast T_T

OpenStudy (usukidoll):

that's like ok distribute a^2 all over now again with the square root la la la la la

OpenStudy (usukidoll):

\[(\frac{ 2a\sqrt{a^2-4b^2}+2a^2-4b^2}{4})(\frac{+ 2a\sqrt{a^2-4b^2}+2a^2-4b^2}{4})\]

OpenStudy (anonymous):

\(a^2+a^2=2a^2\)

OpenStudy (usukidoll):

:/ this is a monsterrrrrr

OpenStudy (usukidoll):

anyone know a mathematica widget or something that can calculate this beast

OpenStudy (anonymous):

Also it might help to do \(2a=\sqrt{4a^2}\)

OpenStudy (usukidoll):

is that legal?

OpenStudy (anonymous):

Well, you don't have to do it.

OpenStudy (anonymous):

You could use this to avoid foiling:\[ (a+b+c)^2 = a^2+b^2+c^2+2ab+2ac+2bc \]

OpenStudy (usukidoll):

o-o

OpenStudy (anonymous):

Another thing is binomial theorem.

OpenStudy (usukidoll):

\[(\frac{ 2a\sqrt{a^2-4b^2}+2a^2-4b^2}{4})\] with a=2a \sqrt{a^2-4b^2} b = 2a^2 and c = 4b^2

OpenStudy (usukidoll):

\[(2a^2)^2 = 4a^4...(4b^2)^2 = 16a^4\]

OpenStudy (anonymous):

\[ (a+b)^4=a^4+4a^3b+6a^2b^2+4ab^3+b^4 \]

OpenStudy (usukidoll):

\[(2a\sqrt{a^2-4b^2})^2 = 4a^2(a^2-4b^2) = 4a^4-16a^2b^2\]

OpenStudy (anonymous):

Is life going well?

OpenStudy (usukidoll):

\[4a^4-16a^2b^2+4a^4+16b^4... 8a^4-16a^2b^2+a16b^4\]

OpenStudy (usukidoll):

\[4a^4-16a^2b^2+4a^4+16b^4... 8a^4-16a^2b^2+16b^4\]

OpenStudy (usukidoll):

like this?!

OpenStudy (usukidoll):

oh wait forgot the 16 from the denominator

OpenStudy (usukidoll):